OpenStudy (anonymous):

How do I factor by grouping 4x-4y+ax-ay?

OpenStudy (anonymous):

4x-4y+ax-ay = 4x+ax-4y-ay = x(4+a)-y(4+a)

OpenStudy (watchmath):

\(4(x-y)+a(x-y)=(4+a)(x-y)\)

OpenStudy (anonymous):

=(4+a)(x-y)

OpenStudy (anonymous):

I need some sort of formula that I can use on the rest of the problems.

OpenStudy (anonymous):

see what the first 2 terms have in common and factor it out, check that you get the same thing when you factor out what the next two terms have in common